A fair die is tossed n times. What is the probability that the sum of the faces showing is n + 2?

Comment: The answer, of course, is not a number but rather an expression involving n).
Answer #1

here total number of outcomes on n dies =6n

number of ways sum of the faces showing is n + 2 =N((n-1) dies shows 1 and one die shows a value 3)+N((n-2) dies shows a number 1 and 2 of n dies shoes numbers 2)

=nC1+nC2 =n+n*(n-1)/2 =n*(2+n-1)/2 =n*(n+1)/2

therefore probability that the sum of the faces showing is n + 2 =n*(n+1)/(2*6n)